Use cross products to determine whether each of pair of ratios forms a proportion:
4/9, 12/27
Please explain how to solve this problem.

Answer by stanbon(75887)

Use cross products to determine whether each of pair of ratios forms a proportion:
4/9, 12/27
------------------
You are checking the following equality:
(4/9)=(12/27)
You do this by cross multiplying as follows:
4*27=108
9*12=108
Since the cross-products are equal the ratios are equal
and they form a proportion.
Cheers,
Stan H.